OpenStudy (anonymous):

Why is the square root of 6 irrational?

OpenStudy (amistre64):

becasue it is not well behaved :)

OpenStudy (anonymous):

because it never ends and doesnt have a pettern

OpenStudy (amistre64):

everytime you try to measure it, it gets fidgety and moves around

OpenStudy (anonymous):

thats hilarious amistre64

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

Let it be rational, and express it as a fraction rt6 = a/b (a/b)(a/b) = 6 a^2/b^2 = 6 a^2 = 6b^2 Since a is even, there is some integer c that is half of a, or in other words: 2c = a. (2c)^2 = (2)(3)b^2 2c^2 = 3b^2 Then b is even and a/b is not in its simplest form.

OpenStudy (anonymous):

(at least I think so)

OpenStudy (anonymous):

whaaat? i did not follow that. luckily, i didnt ask that wuestion

OpenStudy (anonymous):

our initial assumption is that a/b is in its simplest form. I showed that both a and b had to be a multiple of 2.

OpenStudy (anonymous):

merllet15115: Actually, you did. JethroKuan presented a proof of irrationality which shows 'why' \(\sqrt{2}\) is not rational. :)

OpenStudy (anonymous):

Sorry, square root of 6.

OpenStudy (anonymous):

ok. i give up. watevs

OpenStudy (anonymous):

Where did this question come up? What would you consider a good enough reasoning?

OpenStudy (anonymous):

Proof by contradiction should suffice.

OpenStudy (anonymous):

i just havent learned, or atleast i dont think i have, any of that stuff

OpenStudy (anonymous):

And irrational number would by definition mean a number that cannot be presented in the form a/b where a and b are whole numbers. Square root of 6 is irrational since it cannot be written as a fraction. Another feature of irrational numbers is that their decimal forms have an infinite amount of decimals that come up in an unpredictable, non-repeating way. A number with decimal expression 5,67676767676767676767676767676767... is therefore rational since the pair 67 keeps repeating forever. For irrational numbers, it's something complex and non-repeating, for example your \(\sqrt{6}=2.4494897427831780981972840747058913919659474806566701284326...\)